Consider two identical Cournot firms that each have a marginal cost of 20, facing the market inverse demand function: P=120-Q What is the quantity produced for each firm? Round your answer to the nearest 1 decimal places.
Answer #1

Please give ratings it will be appreciable, for any query please comment, Thank you

Solution,

P- 120 - market Inuerse dimand funition where a G + O2 pa 120-01-02 ML = 20 the Equilibrium londition is mr=mc. Total Meenue

0, - 50-0.58, 4 02: 50-015, Q1=50-015 (50-0.50A) + 2. Q, =50-25 t 025 Q -0.25Q = 25 0758, =25 - 25 s 3:33 9.5 Similarly 02=50